A Complete Bibliography of Publications in Computer Networks (Amsterdam, Netherlands: 1999)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

06 March 2014 Version 1.06

Title word cross-reference

T M [122].

1.0 [115]. 1.1 [115]. 1999 [206].

20th [145]. 24th [207].

3 [29]. 36190 [36]. 3D [143].

4 [29].

802.11 [175, 177]. 802.3x [192].

ABR [125]. academic [147]. Access [21, 23, 24, 32, 57, 60, 62, 64, 104, 138, 142, 151, 158, 176]. Accurate [195]. active [42]. activities [15]. adaptable [82]. adaptation [43, 102]. Adaptive [48, 80, 184, 199]. Adding [109]. addresses [165]. addressing [141]. Adjustment [199]. Admission [10, 168]. Advanced [116, 131]. advertising [69, 70, 81]. agent [129, 132–134]. agent-based [132]. Agents [130, 131, 185]. Aggregation [198]. Algorithm [6, 97, 137, 197, 199, 203, 204]. allocation [59, 124]. analysing [152]. Analysis [68, 121, 125, 163, 166]. Analyzing [5]. Annotations [73]. Anticipating [92]. appliance [112]. appliance-independent [112]. Application [42, 46]. Applications [8, 43, 177, 178]. approach [107, 144]. AQuaFWiN [184]. Architectural [21]. Architecture [14, 16, 17, 33, 34, 49, 58, 101, 181, 188]. architectures [18, 41]. Area [28, 139, 176, 205]. ARQ [56]. art [20]. assignment [159]. Asynchronous [174]. ATM [1, 3, 25, 33–39, 56–58, 60– 66, 124, 125, 157, 173, 183, 185, 187]. attack [163]. AU-4 [29]. Authenticating [53]. authorization [105]. Automated [202]. Automatic [84]. automation [120]. available [1, 129]. AWA [57]. AWACS [64]. AXD [35].

back [150]. Backbone [33, 34, 37]. backwards [111]. band [57]. Bandwidth [9, 48, 176, 198, 203]. Bank [168]. base [58]. Based [15, 21, 26, 32, 132, 133, 157, 158, 163, 181, 190, 199, 204]. basic [13]. behavior [102]. Best [197]. better [79]. between [115]. bit [1]. Bluetooth [172]. bounds [10, 65]. Bro [162]. Broadband [11, 20, 23–25, 131, 138, 153]. broadcast [165]. Broadcasts [201]. brokers [68]. browser [86, 144]. Bucket [203]. Buffer [124, 189]. Building [52]. burst [3]. burst-level [3]. Burstiness [203]. bursty [1, 3]. bus [28, 179]. businesses [24].

CAC [174]. cache [87]. caching [77, 79]. call [10]. Canada [206]. candidate [64]. capable [2]. capacity [37, 123]. card [104]. carried [31]. Case [4, 80, 196]. CATV [20]. CATV-HFC [20]. Causal [201]. Causal-Phase [201]. Causality [200]. century [145]. certificates [151]. challenges [85]. Channel [56, 123, 189]. channels [71]. Characteristics [173]. click [69]. clustering [88]. combining [71]. commerce [106]. communication [11, 48, 120, 121]. communications [15, 153]. communities [96]. Comparison [93, 184]. compliant [130]. compression [48]. Computer [207]. concept [146]. conceptual [80]. condensation [136]. Conference [140, 206, 207]. conferences [14]. conferencing [14, 16]. connect [139]. connected [28]. Connections [45, 113, 158, 183]. conscious [44].

Considerations [21, 63]. consistent [77]. Constraint [203]. content [71, 102, 103]. context [134]. Control [10, 16, 38, 125, 130, 157, 168, 192, 200]. Controller [189]. Convergence [27]. cooperative [78, 92]. coordination [16]. coprocessor [52]. core [36]. Correlation [191]. cost [141]. crawling [107]. creating [106]. critical [123]. cross [139]. customer [26]. customization [81, 133]. cyber [96]. cyber-communities [96].

D-ARQ [56]. Data [30, 108, 189, 193]. decisions [43]. Decoupling [203]. DECT [176]. definable [67]. Delay [65, 203]. delay/jitter [65]. Delays [194]. delivery [26, 122]. demand [9]. Design [2, 21, 57, 73, 125, 159, 160]. Designing [39]. detecting [162]. detection [50, 161, 163, 164]. Determining [177]. developing [92]. development [120]. dial [158]. dial-up [158]. differences [115]. Digital [19]. directed [6]. directions [13]. Discovery [6, 84, 107, 185]. disjoint [27]. displays [72]. Distributed [8, 58, 59, 78, 105, 121, 163, 179, 181]. Distribution [194]. division [159]. document [98]. documents [75, 91]. DRS [7]. DSA [56]. dual [28]. Duplex [192]. Dynamic [7, 59, 88, 109, 124, 125, 165, 167, 197].

Easy [179]. Easy-to-Use [179]. Editorial [128, 154]. Effects [135, 192]. Efficiency [9, 68]. Efficient [137, 195]. Eighth [206]. electronic [68, 106, 146]. Embedding [91]. emerging [26, 96]. Enabled [104, 188]. Enabling [43]. End [4, 28, 48, 135, 198]. end-connected [28]. End-to-End [48, 135, 198]. engine [149]. engineering [118]. enhanced [67]. enterprise [38]. Environment [131, 133, 179]. Environments [200]. ESP [37]. Ethernet [141, 192, 195]. European [64, 147].

Evaluating [174]. evaluation [85]. Event [191]. evidence [164]. exhibition [145]. Experiences [169]. Explicit [65]. Extensions [186]. extracting [108]. extreme [39].

fault [7]. FDDI [196]. FETEX [37]. FETEX-150 [37]. Fetuccino [109]. fibre [159]. Filtering [135]. Finding [95]. FIPA [130]. FIPA-compliant [130]. flat [58]. flexible [40]. Flow [31, 157, 189, 192, 198]. Focused [107, 109]. Formal [169, 187]. forms [76]. Formsheets [76]. Framework [80, 163, 184, 202]. Frameworks [169, 184]. free [142]. FSMs [127]. Full [192]. Full-Duplex [192]. function [160]. functions [125]. future [20]. Fuzzy [191].

Gatekeeper [181]. Gecko [114]. general [121]. generation [35, 84]. Genetic [204]. German [145]. GHz [57]. GIF [44]. gigabit [141]. global [101]. graphical [75]. Grouper [88]. Guarantee [203].

H [16]. H-series [16]. H.323 [15, 181]. H.323-Based [181]. Handed [183]. Handed-Off [183]. handoff [61, 63]. hard [10]. hardware [33]. hashing [77]. Health [202]. heterogeneous [87]. heuristics [108]. HFC [20]. HG [157]. HG-RCP [157]. hierarchical [136]. High [41, 52, 57, 153]. high-performance [52]. High-speed [153]. highly [36]. HIPERLAN [174]. HIPERLINK [64]. history [145]. home [133]. homes [24]. host [103]. hosting [101]. hosts [4]. hot [122]. HTTP [113, 115]. HTTP/1.0 [115]. HTTP/1.1 [115]. Hub [205]. Hypercube [190]. hyperlinks [98]. hypermedia [86, 94, 98].

IBM [33, 34]. identification [31]. IEEE [175, 177, 192]. IETF [17]. image [44]. Immune [202]. Impact [18, 29, 30, 62, 167]. Implementation [2, 57, 60, 147].

Implementing [59]. improve [102]. Improved [6, 79, 166]. Improving [9, 72, 100, 110]. incognito [54]. independent [108, 112]. Index [83, 170, 171]. information [66, 89, 97, 110, 136, 151]. infrastructure [92, 105]. Initiators [201]. Integrated [71, 168, 202]. integrated-services [168]. Integrating [94]. Intelligent [130, 205]. inter [61]. inter-switch [61]. interaction [72]. interactive [15, 90]. interconnected [157]. Interconnection [178]. interface [88, 112, 148]. interfering [1]. International [206]. internet [4, 14, 17, 18, 43, 44, 49, 142, 152, 155, 158, 188, 193]. Intra [198]. Intra-Networks [198]. intranets [157]. intruders [162]. Intrusion [50, 161, 163, 164]. intrusion-detection [50]. IP [16, 31, 141, 172]. ISDN [16, 155]. ITU [15, 16]. ITU-T [15, 16].

Joint [56].

Key [115]. knowledge [91]. KPS [97].

LAN [32, 58, 175, 204]. language

[74–76, 112]. LANs [173, 177, 192]. Large [14, 37, 38, 163]. latency [5]. Layer [174]. LCN’99 [207]. Leading [172]. Leaky [203]. Learning [108]. LeMO [145]. lens [100]. Level [3, 10, 42, 46, 173]. Licence [142]. Licence-free [142]. Lightwave [137, 205]. lightweight [70]. limited [121]. lines [19]. Link [55, 99, 100, 189]. links [6]. LMDS [24]. Local [22, 138, 139, 176, 205, 207]. loop [22]. low [141]. low-cost [141]. Lowell [207]. Lowering [55].

MAC [56, 141, 173]. magazine [146]. MainStreetXpress [36]. Management [62, 66, 89, 144, 165, 176, 189, 191, 202]. Managing [46, 113, 188]. mapping [29]. marketplace [68]. Massachusetts [207]. Maximum [122]. May [206]. Measured [175]. Measurement [158, 182]. Measurement-based [158]. Measuring [5, 83]. mechanism [32, 87, 121, 157]. mechanisms [13, 86]. Media [94, 200]. medical [104]. MEHARI [152]. merging [141]. message [123, 137]. metadata [84, 93]. metropolitan [28]. middleware [132]. minimum [123]. mining [97]. Mirror [103]. MMDS [24]. Mobile [47, 129, 131, 133, 151, 185, 186]. Mobility [54, 66, 132, 134, 172]. Model [178, 191]. modeled [127]. Modelling [158, 169, 176]. Models [56, 193]. movies [143]. movies-real-time [143]. MSC [116]. MSC’96 [117]. Multi [138, 189, 201]. multi-access [138]. Multi-Channel [189]. Multi-Initiators [201]. multiaccess [159]. Multicast [5, 8, 186, 197]. multicasting [11]. Multimedia [10, 14–16, 71, 82, 92, 181, 184]. My [148, 149].

nature [51]. navigational [110]. Network [7, 26, 44, 130, 135, 147, 160, 162, 167, 176, 180, 185, 191, 196, 200, 202, 204]. network-based [26]. Network-conscious [44]. Networked [178]. networking [42]. Networks [3, 6, 10, 11, 13, 15, 16, 20, 21, 28, 31, 47, 61, 62, 66, 136–139, 141, 153, 156, 157, 159, 165, 168, 183, 184, 186, 190, 194, 198, 205, 207]. Neural [204]. newly [26]. newspaper [71]. NFS [114]. node [38]. noisy [123]. Novel [204].

object [169]. object-oriented [169]. objects [143]. Observers [187]. October [207]. Off [183]. open [86, 94]. Operational [117]. Optical [21, 159, 160]. Optimal [190]. Optimization [61, 183]. order [45]. Ordering [201]. organization [89]. oriented [169]. other [69, 184]. out-of-sequence [30]. Outdoor [177]. overhead [55]. Overlapped [201]. overload [102].

Packet [13, 15, 173, 199]. packet-based [15]. page [108]. page-independent [108]. pages [95, 108]. pairs [103]. pan [147]. pan-European [147]. paper [140]. paradigm [26]. parameters [127]. partial [45]. Path [61, 197]. patient [104]. patterns [110]. pay [69]. pay-per-click [69]. PCN [63]. IP [194]. jitter [65]. MSC [118]. per-VC [38]. Performance [1, 29, 30, 41, 52, 56, 118, 135, 159, 173–175]. Periodic [183]. Periods [199]. persistent [113]. personal [71, 134, 149]. personalizable [89]. perspective [4, 17]. PFQ [203]. Phase [201]. PKI [51]. placement [160]. platforms [129]. Playout [199]. pointer [29]. potential [64]. PowerBookmarks [89]. Prediction [204]. preestablished [63]. premises [26]. presentations [82]. preview [99]. printed [98]. priority [3, 168]. Proceedings [206, 207]. processing [30]. processors [29]. Programmable [47, 52]. Programming [180]. Propagation [177]. Protocol [8, 41, 56, 61, 159, 173, 174, 201, 205]. Protocols [5, 17, 63, 120, 123, 127, 187]. Prototype [57]. Providing [151]. proxy [86, 114]. public [53]. publishing [71].

QoS [2, 43, 182, 184, 188, 203]. QoS-Enabled [188]. Quality [10, 13, 46, 83]. Quality-of-service [13]. query [74]. querying [75, 87]. questions [148]. queue [125]. queueing [135].

RaDaR [101]. Radio [22, 24, 67, 177]. random [83]. Range [177]. rate [1, 157]. rate-based [157]. RCP [157]. RDF [84]. real [8, 65, 130, 143, 156, 162]. real-time [8, 65, 130, 156, 162]. realistic [56]. Reality [143, 178]. Reasoning [191]. recommendations [15, 16]. record [104]. records [106]. recursive [126]. Regulation [32, 198]. related [15, 95]. Reliability [4, 39]. reliable [36]. replicated [103]. representation [93]. rerouting [63]. research [147]. resource [59, 66, 84, 107]. Resources [79, 179]. responses [79]. restructuring [75]. Results [85, 88]. retailer [134]. Retaining [98]. ring [159]. robin [126, 198]. Round [126, 194]. Round-Trip [194]. Route [183]. router [2]. Routing [6–8, 27, 55, 190]. RRR [126]. runtime [90].

s [128, 140, 154]. satellite [23]. SBPT [197]. Scalability [198]. scalable [36, 101]. scale [38, 163]. Scenario [196]. scheduler [126]. Scheduling [11, 121, 137, 156, 168]. schemas [93]. Scheme [3, 176, 200]. schemes [69, 125]. SDH [29]. SDL [116, 118]. SDL/ MSC [118]. search [85, 88, 109, 148, 149]. Secure [52, 70, 151, 186]. Security [49, 55, 69, 202]. selection [134, 167]. Semantic [87]. Semantics [117]. sequence [30]. Series [16, 193]. server [79, 102, 167]. ServerNet [122]. servers [9, 78, 151]. Service [10, 13, 46, 101, 119, 131, 133, 134, 169]. Services [26, 36, 90, 144, 152, 168, 172, 180]. Session [165]. sessions [11]. sharing [89]. Shortest [197]. SHRs [166]. Signaling [187]. signalling [61]. Silence [199]. simple [105]. Simulation [195]. sites [80]. Skew [182]. small [72]. smart [104]. SmartARP [141]. SMIL [92]. Software [34, 67, 202]. solution [132]. SONET [166]. sources [1]. spanning [27]. spatial [139]. specific [107]. Specification [187]. Specify [187]. speed [57, 153]. spots [122]. SR [56]. SR/D [56]. SR/D-ARQ [56]. standard [64]. standardization [15]. standards [129]. state [20, 55]. state-of-the-art [20]. stations [58]. strategy [163]. streaming [73, 143]. structuring [86]. studies [73]. Study [4, 80, 103, 182]. subscriber [19]. subscription [133]. Super [190]. supercomputer [153]. support [11, 86, 109]. Surfing [111]. survivability [166]. switch [2, 32–34, 36, 37, 39, 40, 61, 124, 125]. switch-based [32]. switch-router [2]. switched [31]. switching [35, 58]. Synchronization [182, 200]. Synchronous [196]. System [7, 35, 57, 64, 67, 73, 89, 90, 121, 152, 162, 202, 204]. systems [22, 25, 50, 110, 118, 135, 164].

T [15, 16]. Talking [150]. Talkspurt [199]. taxonomy [50]. TCP [1, 113, 194]. TCP/ IP [194]. technique [138]. techniques [81]. technologies [142]. technology [134]. telecommunication [18, 169]. Telephony [12, 17, 181]. Temporal [94, 191]. temporary [151]. TEN [147]. TEN-155 [147]. terabit [40]. terminal [60, 132]. terminals [53]. Test [120]. Testing [127]. Time [8, 65, 122, 123, 130, 143, 156, 162, 193]. time-critical [123]. Timestamp [156]. Timestamp-Scheduling [156]. timing [127]. Token [168]. tolerance [7]. TOMTEN [46]. topic [107]. topic-specific [107]. Topics [116]. topologies [122]. Topology [6, 28, 136, 190]. Toronto [206]. Totem [5]. Traffic [1, 3, 31, 38, 56, 62, 65, 130, 174, 193, 196, 203]. transaction [106]. transit [40]. Transmission [30, 44, 123, 196]. Transport [45]. travel [54]. traversal [99]. Trawling [96]. Tree [63, 197]. Trees [27, 197]. trends [20]. Trip [194]. TU [29]. TU-3 [29]. TV [71]. Two [20]. Two-way [20].

ubiquitous [104]. UIML [112]. UMTS [133]. understanding [79]. uniform [29]. Unintrusive [81]. Universal [156]. Untraceable [54]. USA [207]. usability [100]. usage [73]. Use [134, 152, 179]. useable [51]. User [4, 82, 112, 132]. Using [16, 57, 83, 106, 133, 139, 169, 185, 187].

Validate [187]. validation [119]. variable [48]. variable-bandwidth [48]. VC [38]. Very [14]. via [23]. video [9, 45, 73, 93]. video-on-demand [9]. Virtual [133, 143, 145, 178, 179]. Visual [99]. Voice [199]. VPNs [188].

walks [83]. wander [29]. WATM [59]. wavelength [139, 159]. Way [20, 172]. WDAI [105]. WDM [21, 137, 190]. WDM-based [21]. WDMA [160]. Web [69, 70, 72, 73, 77–83, 85–92, 94–97, 99– 105, 107, 108, 110, 111, 114, 144, 206]. Web-enabled [104]. Webvise [86]. Wide [82, 86, 94, 95, 99, 105, 206]. Wireless [10, 22, 25, 56–58, 60– 66, 138, 142, 173, 175, 176, 183, 184, 204]. within [131]. wmATM [67]. World [82, 86, 94, 95, 99, 105, 206]. wormhole [6]. Worst [196]. Worst-Case [196]. WWW [150].

XFDL [106]. XML [74–76, 106, 112]. XML-GL [75].

year [12].

References

Hong:1999:PAA

[1] Duke P. Hong and Tatsuya Suda. Performance of ATM available bit rate for bursty TCP sources and interfering traffic. Computer Networks (Amsterdam, Netherlands: 1999), 31(1–2):7–18, January 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/1-
2/2032.pdf.

Basturk:1999:DIQ

[2] E. Basturk, A. Birman, G. Delp, R. Guérin, R. Haas, S. Kamat, D. Kandlur, P. Pan, D. Pendarakis, V. Peris, R. Rajan, D. Saha, and D. Williams.

Design and implementation of a QoS capable switch-router. Computer Networks (Amsterdam, Netherlands: 1999), 31(1–2):19–32, January 14, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/1-2/2033.pdf.

Fernandez:1999:BLP

[3] Jose Roberto Fernandez and Matt W. Mutka. A burst-level priority scheme for bursty traffic in ATM networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(1–2):33–45, January 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/1-
2/2034.pdf.

Kalyanakrishnan:1999:RIH

[4] M. Kalyanakrishnan, R. K. Iyer, and J. U. Patel. Reliability of Internet hosts: a case study from the end user’s perspective. Computer Networks (Amsterdam, Netherlands: 1999), 31 (1–2):47–57, January 14, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/1-2/2035.pdf.

Thomopoulos:1999:AML

[5] E. Thomopoulos, L. E. Moser, and P. M. Melliar-Smith. Analyzing and measuring the latency of the Totem multicast protocols. Computer Networks (Amsterdam, Netherlands: 1999), 31 (1–2):59–78, January 14, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/1-2/2036.pdf.

Huang:1999:ITD

[6] Ying-Yi Huang and Mart L. Molle. An improved topology discovery algorithm for networks with wormhole routing and directed links. Computer Networks (Amsterdam, Netherlands: 1999), 31(1–2):79–88, January 14, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/1-2/2037.pdf.

Chowdhury:1999:DRS

[7] A. Chowdhury, O. Frieder, E. Burger, D. Grossman, and K. Makki. Dynamic Routing System (DRS): fault tolerance in network routing. Computer Networks (Amsterdam, Netherlands: 1999), 31 (1–2):89–99, January 14, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/1-2/2038.pdf.

Jia:1999:DMR

[8] Xiaohua Jia, Yanchun Zhang, Niki Pissinou, and Kia Makki. A distributed multicast routing protocol for real-time multicast applications. Computer Networks (Amsterdam, Netherlands: 1999), 31(1–2):101–110, January 14, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/1-2/2040.pdf.

Carter:1999:IBE

[9] Steven W. Carter and Darrell D. E. Long. Improving bandwidth efficiency of video-on-demand servers. Computer Networks (Amsterdam, Netherlands: 1999), 31(1–2):111–123, January 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/1-
2/2039.pdf.

Misic:1999:ACW

[10] Jelena Mišić, Samuel T. Chanson, and Frederick S. Lai. Admission control for wireless multimedia networks with hard call level quality of service bounds. Computer Networks (Amsterdam, Netherlands: 1999), 31(1–2):125–140, January 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/1-
2/2041.pdf.

Sheta:1999:SSM

[11] Khalid H. Sheta and Mukesh Singhal. Scheduling support for multicasting sessions in broadband communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 31 (1–2):141–150, January 14, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/1-2/2042.pdf.

Rinde:1999:TY

[12] Joseph Rinde. Telephony in the year 2005. Computer Networks (Amsterdam, Netherlands: 1999), 31(3): 157–168, February 11, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/3/2072.pdf.

Guerin:1999:QSP

[13] R. Guérin and V. Peris. Quality-of-service in packet networks: basic mechanisms and directions. Computer Networks (Amsterdam, Netherlands: 1999), 31(3):169–189, February 11, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/3/2067.pdf.

Handley:1999:VLC

[14] M. Handley, J. Crowcroft, C. Bormann, and J. Ott. Very large conferences on the Internet: the Internet multimedia conferencing architecture. Computer Networks (Amsterdam, Netherlands: 1999), 31(3):191–204, February 11, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/3/2068.pdf.

Toga:1999:ISA

[15] James Toga and Jörg Ott. ITU-T standardization activities for interactive multimedia communications on packetbased networks: H.323 and related recommendations. Computer Networks (Amsterdam, Netherlands: 1999), 31 (3):205–223, February 11, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/3/2073.pdf.

Reid:1999:MCI

[16] Mark Reid. Multimedia conferencing over ISDN and IP networks using ITU-T H-series recommendations: architecture, control and coordination. Computer Networks (Amsterdam, Netherlands: 1999), 31(3):225–235, February 11, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/
                                        store/comnet/sub/1999/31/3/2070.
                                        pdf.

Schulzrinne:1999:ITA

[17] Henning Schulzrinne and Jonathan Rosenberg. Internet Telephony: architecture and protocols — an IETF perspective. Computer Networks (Amsterdam, Netherlands: 1999), 31(3): 237–255, February 11, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/3/2071.pdf.

Hubaux:1999:IIT

[18] Jean-Pierre Hubaux, Constant Gbaguidi, Shawn Koppenhoefer, and Jean-Yves Le Boudec. The impact of the Internet on telecommunication architectures. Computer Networks (Amsterdam, Netherlands: 1999), 31(3):257–273, February 11, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/
                                        store/comnet/sub/1999/31/3/2069.
                                        pdf.

Cioffi:1999:DSL

[19] J. Cioffi, P. Silverman, and T. Starr. Digital subscriber lines. Computer Networks (Amsterdam, Netherlands: 1999), 31(4):283–311, February 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/4/2079.pdf.

Perkins:1999:TWB

[20] Stephen Perkins and Alan Gatherer. Two-way broadband CATV-HFC networks: state-of-the-art and future trends. Computer Networks (Amsterdam, Netherlands: 1999), 31(4): 313–326, February 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/4/2078.pdf.

Modiano:1999:ACD

[21] Eytan Modiano and Richard Barry. Architectural considerations in the design of WDM-based optical access networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(4): 327–341, February 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/4/2075.pdf.

Cotanis:1999:WLL

[22] Nicolae Cotanis and Bijan Jabbari. Wireless local loop radio systems. Computer Networks (Amsterdam, Netherlands: 1999), 31(4):343–352, February 25, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/4/2077.
pdf.

Hadjitheodosiou:1999:BAS

[23] M. H. Hadjitheodosiou, A. Ephremides, and D. Friedman. Broadband access via satellite. Computer Networks (Amsterdam, Netherlands: 1999), 31 (4):353–378, February 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/4/2080.pdf.

Sari:1999:BRA

[24] Hikmet Sari. Broadband radio access to homes and businesses: MMDS and LMDS. Computer Networks (Amsterdam, Netherlands: 1999), 31(4): 379–393, February 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/4/2076.pdf.

Ayanoglu:1999:WBA

[25] Ender Ayanoglu. Wireless broadband and ATM systems. Computer Networks (Amsterdam, Netherlands: 1999), 31 (4):395–409, February 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/4/2082.pdf.

Waring:1999:NEC

[26] David L. Waring, Kenneth J. Kerpez, and Steven G. Ungar. A newly emerging customer premises paradigm for delivery of network-based services. Computer Networks (Amsterdam, Netherlands: 1999), 31(4):411–424, February 25, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/
                                        store/comnet/sub/1999/31/4/2074.
                                        pdf.

Yener:1999:CRD

[27] Bülent Yener, Yoram Ofek, and Moti Yung. Convergence routing on disjoint spanning trees. Computer Networks (Amsterdam, Netherlands: 1999), 31(5):429–443, March 11, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/5/2065.pdf.

Ramapriya:1999:ECD

[28] M. R. Ramapriya, S. Beotra, M. Mukherjee, R. V. Raja Kumar, and P. Dhar. An end-connected dual bus topology for metropolitan area networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(5):445–456, March 11, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/5/2083.
pdf.

Sholander:1999:IMW

[29] Peter E. Sholander and Henry L. Owen. The impact of mapping wander on the performance of SDH AU-4 and TU-3 uniform pointer processors. Computer Networks (Amsterdam, Netherlands: 1999), 31(5):457–473, March 11, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/5/2084.
pdf.

Diot:1999:ISP

[30] Christophe Diot and François Gagnon. Impact of out-of-sequence processing on the performance of data transmission. Computer Networks (Amsterdam, Netherlands: 1999), 31(5): 475–492, March 11, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/5/2093.pdf.

Copeland:1999:IFI

[31] John A. Copeland, Randal Abler, and Keith L. Bernhardt. IP flow identification for IP traffic carried over switched networks. Computer Networks (Amsterdam, Netherlands: 1999), 31 (5):493–504, March 11, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/5/2096.pdf.

Ofek:1999:ARM

[32] Yoram Ofek and Moti Yung. Access regulation mechanism for switchbased LAN. Computer Networks (Amsterdam, Netherlands: 1999), 31 (5):505–518, March 11, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/5/2097.pdf.

Alaiwan:1999:IABa

[33] Haissam Alaiwan. IBM 8265 ATM Backbone Switch hardware architecture. Computer Networks (Amsterdam, Netherlands: 1999), 31(6): 527–539, March 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/6/2086.pdf.

Alaiwan:1999:IABb

[34] Haissam Alaiwan. IBM 8265 ATM Backbone Switch software architecture. Computer Networks (Amsterdam, Netherlands: 1999), 31(6):541–558, March 25, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/

store/comnet/sub/1999/31/6/2087.
pdf.

Blau:1999:ANG

[35] Staffan Blau, Jan Rooth, Jörgen Axell, Fiffi Hellstrand, Magnus Buhrgard, Tommy Westin, and Göran Wicklund. AXD 301: A new generation ATM switching system. Computer Networks (Amsterdam, Netherlands: 1999), 31 (6):559–582, March 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/6/2088.pdf.

Rathgeb:1999:MSH

[36] Erwin P. Rathgeb. The MainStreetXpress 36190: a scalable and highly reliable ATM core services switch. Computer Networks (Amsterdam, Netherlands: 1999), 31(6):583–601, March 25, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/6/2089.
pdf.

Soumiya:1999:LCA

[37] Toshio Soumiya, Koji Nakamichi, Satoshi Kakuma, Takashi Hatano, and Akira Hakata. The large capacity ATM backbone switch “FETEX-150 ESP”. Computer Networks (Amsterdam, Netherlands: 1999), 31(6): 603–615, March 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/6/2090.pdf.

Endo:1999:LSE

[38] Noboru Endo, Takeshi Aimoto, Takeki Yazaki, and Tomihisa Nishijima. Large scale enterprise ATM node with per-VC traffic control. Computer Networks (Amsterdam, Netherlands: 1999), 31 (6):617–625, March 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/6/2091.pdf.

Hermelee:1999:DAS

[39] Alan Hermelee and Gary McElvany. Designing an ATM switch for extreme reliability. Computer Networks (Amsterdam, Netherlands: 1999), 31 (6):627–637, March 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/6/2092.pdf.

Osstyn:1999:FTT

[40] E. Osstyn. A flexible terabit transit switch. Computer Networks (Amsterdam, Netherlands: 1999), 31(6): 639–648, March 25, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/6/2095.pdf.

Crowcroft:1999:HPP

[41] Jon Crowcroft and Cristophe Diot. High Performance Protocol Architectures. Computer Networks (Amsterdam, Netherlands: 1999), 31(7):653–654, April 8, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/7/
2098.pdf.

Fry:1999:ALA

[42] Michael Fry and Atanu Ghosh. Application level active networking. Computer Networks (Amsterdam, Netherlands: 1999), 31(7):655–667, April 8, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/7/2099.
pdf.

Bhatti:1999:EQA

[43] Saleem N. Bhatti and Graham Knight. Enabling QoS adaptation decisions for Internet applications. Computer Networks (Amsterdam, Netherlands: 1999), 31(7):669–692, April 8, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/7/2100.pdf.

Amer:1999:NCG

[44] Paul D. Amer, Sami Iren, Gul E. Sezen, Phillip T. Conrad, Mason Taube, and Armando Caro. Network-conscious GIF image transmission over the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 31(7):693–708, April 8, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/7/
2101.pdf.

Rojas-Cardenas:1999:TVP

[45] Luis Rojas-Cárdenas, Emmanuel Chaput, Laurent Dairaine, Patrick Sénac, and Michel Diaz. Transport of video over partial order connections. Computer Networks (Amsterdam, Netherlands: 1999), 31(7):709–725, April 8, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/7/2102.
pdf.

Silva:1999:MAL

[46] R. De Silva, B. Landfeldt, S. Ardon, A. Seneviratne, and Christophe Diot. Managing application level quality of service through TOMTEN. Computer Networks (Amsterdam, Netherlands: 1999), 31(7):727–739, April 8, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/7/2103.pdf.

Campbell:1999:PMN

[47] Andrew T. Campbell, Michael E. Kounavis, and Raymond R.-F. Liao. Programmable mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(7):741–765, April 8, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/7/2104.
pdf.

Knutsson:1999:AEE

[48] Björn Knutsson and Mats Björkman. Adaptive end-to-end compression for variable-bandwidth communication. Computer Networks (Amsterdam, Netherlands: 1999), 31(7):767–779, April 8, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/7/2105.
pdf.

Molva:1999:ISA

[49] Refik Molva. Internet security architecture. Computer Networks (Amsterdam, Netherlands: 1999), 31(8): 787–804, April 23, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/8/2121.pdf.

Debar:1999:TTI

[50] Hervé Debar, Marc Dacier, and Andreas Wespi. Towards a taxonomy of intrusion-detection systems. Computer Networks (Amsterdam, Netherlands: 1999), 31(8):805–822, April 23, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/8/2122.
pdf.

Ellison:1999:NUP

[51] Carl M. Ellison. The nature of a useable PKI. Computer Networks (Amsterdam, Netherlands: 1999), 31 (8):823–830, April 23, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/8/2123.pdf.

Smith:1999:BHP

[52] Sean W. Smith and Steve Weingart. Building a high-performance, programmable secure coprocessor. Computer Networks (Amsterdam, Netherlands: 1999), 31(8):831–860, April 23, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/8/2124.
pdf.

Asokan:1999:APT

[53] N. Asokan, Hervé Debar, Michael Steiner, and Michael Waidner. Authenticating public terminals. Computer Networks (Amsterdam, Netherlands: 1999), 31(8):861–870, April 23, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/
                                        store/comnet/sub/1999/31/8/2125.
                                        pdf.

Ateniese:1999:UMH

[54] Giuseppe Ateniese, Amir Herzberg, Hugo Krawczyk, and Gene Tsudik. Untraceable mobility or how to travel incognito. Computer Networks (Amsterdam, Netherlands: 1999), 31(8): 871–884, April 23, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/8/2126.pdf.

Hauser:1999:LSO

[55] Ralf Hauser, Tony Przygienda, and Gene Tsudik. Lowering security overhead in link state routing. Computer Networks (Amsterdam, Netherlands: 1999), 31(8):885–894, April 23, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/
                                        store/comnet/sub/1999/31/8/2127.
                                        pdf.

Petras:1999:JPD

[56] Dietmar Petras and Ulrich Vornefeld. Joint performance of DSA++ MAC protocol and SR/D-ARQ protocol for wireless ATM under realistic traffic and channel models. Computer Networks (Amsterdam, Netherlands: 1999), 31 (9–10):903–918, May 7, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/9-10/2107.pdf.

Manabe:1999:PAD

[57] Toshihiro Manabe, Takatoshi Sugiyama, Atsushi Ohta, Fusao Nuno, Hiroyuki Suzuki, Shuta Uwano, Satoru Aikawa, and Masahiro Umehira. Prototype AWA: design and implementation of a high speed ATM wireless access system using 25 GHz band. Computer Networks (Amsterdam, Netherlands: 1999), 31(9–10):919–930, May 7, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/9-10/2108.pdf.

Du:1999:FWA

[58] Y. Du, C. Herrmann, K. P. May, and David Evans. A flat wireless ATM LAN architecture with distributed ATM switching in base stations. Computer Networks (Amsterdam, Netherlands: 1999), 31(9–10):931–942, May 7, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/9-10/2109.pdf.

Sinner:1999:IDD

[59] Christian Sinner and Michael Wolf. Implementing distributed and dynamic resource allocation in WATM. Computer Networks (Amsterdam, Netherlands: 1999), 31(9–10):943–957, May 7, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/

                                        tree/store/comnet/sub/1999/31/9-
                                        10/2110.pdf.

Ala-Laurila:1999:IWA

[60] Juha Ala-Laurila and Jussi Lemilainen. Implementation of the wireless ATM access terminal. Computer Networks (Amsterdam, Netherlands: 1999), 31 (9–10):959–973, May 7, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/9-10/2111.pdf.

Wong:1999:POS

[61] W. S. Vincent Wong and Victor C. M. Leung. A path optimization signalling protocol for inter-switch handoff in wireless ATM networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(9–10):975–984, May 7, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/9-10/2112.pdf.

Sigle:1999:IWA

[62] Rolf Sigle and Thomas Renger. Impact of wireless access on traffic management in ATM networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(9–10):985–997, May 7, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/9-10/2113.pdf.

Jiang:1999:CPT

[63] Jianping Jiang, Ten-Hwang Lai, and Min-Te Sun. Considerations on preestablished tree rerouting handoff protocols for wireless ATM PCN. Computer Networks (Amsterdam, Netherlands: 1999), 31(9–10):999–1009, May 7, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/9-
10/2114.pdf.

Evci:1999:AAW

[64] C. Evci, A. De Hoz, R. Rheinschmitt, M. Araki, M. Beach, A. Nix, P. Hafezi, Y. Sun, S. Barberis, E. Gaiani, B. Melis, G. Romano, V. Palestini, M. Tolonen, and H. Hakalahti. AWACS: ATM Wireless Access System — potential candidate for European HIPERLINK standard. Computer Networks (Amsterdam, Netherlands: 1999), 31(9– 10):1011–1028, May 7, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/9-10/2115.pdf.

Pocher:1999:EDJ

[65] Hervé Le Pocher, Victor C. M. Leung, and Donald Gillies. Explicit delay/jitter bounds for real-time traffic over wireless ATM. Computer Networks (Amsterdam, Netherlands: 1999), 31 (9–10):1029–1048, May 7, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/9-10/2116.pdf.

Shen:1999:MIR

[66] Xuemin Shen and Jon W. Mark. Mobility information for resource management in wireless ATM networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(9–10):1049–1062, May 7, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/31/9-
                                        10/2117.pdf.

Hu:1999:WES

[67] Janny Hu. wmATM enhanced software definable radio system. Computer Networks (Amsterdam, Netherlands: 1999), 31(9–10):1063–1071, May 7, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/store/
                                        comnet/sub/1999/31/9-10/2118.pdf.

Almeida:1999:EAB

[68] Virgílio A. F. Almeida, Wagner Meira, Jr., Victor F. Ribeiro, and Nivio Ziviani. Efficiency analysis of brokers in the electronic marketplace. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1079–1090, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/31/11-
                                        16/2177.pdf.

Anupam:1999:SPP

[69] Vinod Anupam, Alain Mayer, Kobbi Nissim, Benny Pinkas, and Michael K. Reiter. On the security of payper-click and other Web advertising schemes. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1091–1100, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/
                                        store/comnet/sub/1999/31/11-16/
                                        2149.pdf.

Jakobsson:1999:SLA

[70] Markus Jakobsson, Philip D. MacKenzie, and Julien P. Stern. Secure and lightweight advertising on the Web. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1101–1109, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/11-16/
2167.pdf.

Sodergaard:1999:IMP

[71] Caj Södergård, Matti Aaltonen, Sari Hagman, Mikko Hiirsalmi, Timo Järvinen, Eija Kaasinen, Timo Kinnunen, Juha Kolari, Jouko Kunnas, and Antti Tammela. Integrated multimedia publishing: combining TV and newspaper content on personal channels. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1111–1128, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2156.pdf.

Jones:1999:IWI

[72] Matt Jones, Gary Marsden, Norliza Mohd-Nasir, Kevin Boone, and George Buchanan. Improving Web interaction on small displays. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1129–1137, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2139.pdf.

Bargeron:1999:ASV

[73] David Bargeron, Anoop Gupta, Jonathan Grudin, and Elizabeth Sanocki. Annotations for streaming video on the Web: system design and usage studies. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1139–1153, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/31/11-
                                        16/2183.pdf.

Deutsch:1999:QLX

[74] Alin Deutsch, Mary Fernandez, Daniela Florescu, Alon Levy, and Dan Suciu. A query language for XML. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1155–1169, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/31/11-
                                        16/2146.pdf.

Ceri:1999:XGL

[75] Stefano Ceri, Sara Comai, Ernesto Damiani, Piero Fraternali, Stefano Paraboschi, and Letizia Tanca. XML-GL: a graphical language for querying and restructuring XML documents. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1171–1187, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/31/11-
                                        16/2140.pdf.

Kristensen:1999:FXF

[76] Anders Kristensen. Formsheets and the XML forms language. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1189–1201, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2155.pdf.

Karger:1999:WCC

[77] David Karger, Alex Sherman, Andy Berkheimer, Bill Bogstad, Rizwan Dhanidina, Ken Iwamoto, Brian Kim, Luke Matkins, and Yoav Yerushalmi. Web caching with consistent hashing. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16): 1203–1213, May 17, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/11-16/
2181.pdf.

Baker:1999:DCW

[78] Scott M. Baker and Bongki Moon. Distributed cooperative Web servers. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1215–1229, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2137.pdf.

Wills:1999:TBU

[79] Craig E. Wills and Mikhail Mikhailov. Towards a better understanding of Web resources and server responses for improved caching. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1231–1243, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/31/11-
                                        16/2163.pdf.

Perkowitz:1999:TAW

[80] Mike Perkowitz and Oren Etzioni. Towards adaptive Web sites: conceptual framework and case study. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1245–1258, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/31/11-
                                        16/2143.pdf.

Langheinrich:1999:UCT

[81] Marc Langheinrich, Atsuyoshi Nakamura, Naoki Abe, Tomonari Kamba, and Yoshiyuki Koseki. Unintrusive customization techniques for Web advertising. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1259–1272, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/
                                        store/comnet/sub/1999/31/11-16/
                                        2159.pdf.

Rousseau:1999:UAM

[82] Franck Rousseau, J. Antonio García-Macías, José Valdeni de Lima, and Andrzej Duda. User adaptable multimedia presentations for the World Wide Web. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1273–1290, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/
                                        store/comnet/sub/1999/31/11-16/
                                        2162.pdf.

Henzinger:1999:MIQ

[83] Monika R. Henzinger, Allan Heydon, Michael Mitzenmacher, and Marc Najork. Measuring index quality using random walks on the Web. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1291–1303, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2142.pdf.

Jenkins:1999:ARM

[84] Charlotte Jenkins, Mike Jackson, Peter Burden, and Jon Wallis. Automatic RDF metadata generation for resource discovery. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1305–1320, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2138.pdf.

Hawking:1999:RCW

[85] David Hawking, Nick Craswell, Paul Thistlewaite, and Donna Harman. Results and challenges in Web search evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1321–1330, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/11-16/
2150.pdf.

Gronbaek:1999:WBP

[86] Kaj Grønbæk, Lennert Sloth, and Peter Ørbæk. Webvise: browser and proxy support for open hypermedia structuring mechanisms on the World Wide Web. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1331–1345, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2172.pdf.

Chidlovskii:1999:SCM

[87] Boris Chidlovskii, Claudia Roncancio, and Marie-Luise Schneider. Semantic cache mechanism for heterogeneous Web querying. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1347–1360, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2161.pdf.

Zamir:1999:GDC

[88] Oren Zamir and Oren Etzioni. Grouper: a dynamic clustering interface to Web search results. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1361–1374, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2180.pdf.

Li:1999:PSP

[89] Wen-Syan Li, Quoc Vu, Divakant Agrawal, Yoshinori Hara, and Hajime Takano. PowerBookmarks: a system for personalizable Web information organization, sharing, and management. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1375–1389, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2158.pdf.

Brabrand:1999:RSI

[90] Claus Brabrand, Anders Møller, Anders Sandholm, and Michael I. Schwartzbach. A runtime system for interactive Web services. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1391–1401, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/11-16/
2151.pdf.

Martin:1999:EKW

[91] Philippe Martin and Peter Eklund. Embedding knowledge in Web documents. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1403–1419, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/11-16/
2145.pdf.

Rutledge:1999:ASD

[92] Lloyd Rutledge, Jacco van Ossenbruggen, Lynda Hardman, and Dick C. A. Bulterman. Anticipating SMIL 2.0: the developing cooperative infrastructure for multimedia on the Web. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1421–1430, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/
                                        store/comnet/sub/1999/31/11-16/
                                        2152.pdf.

Hunter:1999:CSV

[93] Jane Hunter and Liz Armstrong. A comparison of schemas for video metadata representation. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1431–1451, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/31/11-
                                        16/2179.pdf.

Bouvin:1999:ITM

[94] Niels Olof Bouvin and René Schade. Integrating temporal media and open hypermedia on the World Wide Web. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1453–1465, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/
                                        store/comnet/sub/1999/31/11-16/
                                        2164.pdf.

Dean:1999:FRP

[95] Jeffrey Dean and Monika R. Henzinger. Finding related pages in the World Wide Web. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1467–1479, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/31/11-
                                        16/2148.pdf.

Kumar:1999:TWE

[96] Ravi Kumar, Prabhakar Raghavan, Sridhar Rajagopalan, and Andrew Tomkins. Trawling the Web for emerging cyber-communities. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1481–1493, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2166.pdf.

Guan:1999:KWI

[97] Tao Guan and Kam-Fai Wong. KPS: a Web information mining algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1495–1507, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/11-16/
2174.pdf.

Wan:1999:RHP

[98] Ernest Wan, Philip Robertson, John Brook, Stephen Bruce, and Kristine Armitage. Retaining hyperlinks in printed hypermedia document. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1509–1524, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2175.pdf.

Kopetzky:1999:VPL

[99] Theodorich Kopetzky and Max Mühlhäuser. Visual preview for link traversal on the World Wide Web. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1525–1532, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2176.pdf.

Stanyer:1999:IWU

[100] Dominic Stanyer and Rob Procter. Improving Web usability with the link lens. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1533–1544, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/11-16/
2160.pdf.

Rabinovich:1999:RSA

[101] Michael Rabinovich and Amit Aggarwal. RaDaR: a scalable architecture for a global Web hosting service. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1545–1561, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2169.pdf.

Abdelzaher:1999:WCA

[102] Tarek F. Abdelzaher and Nina Bhatti. Web content adaptation to improve server overload behavior. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1563–1577, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2157.pdf.

Bharat:1999:MMW

[103] Krishna Bharat and Andrei Broder.

Mirror, mirror on the Web: a study of host pairs with replicated content. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1579–1590, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2147.pdf.

Chan:1999:WES

[104] Alvin T. S. Chan. Web-enabled smart card for ubiquitous access of patient’s medical record. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1591–1598, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2182.pdf.

Kahan:1999:WSW

[105] José Kahan. WDAI: a simple World Wide Web distributed authorization infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1599–1609, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2153.pdf.

Blair:1999:XCE

[106] Barclay Blair and John Boyer. XFDL:

creating electronic commerce transaction records using XML. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1611–1622, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2154.pdf.

Chakrabarti:1999:FCN

[107] Soumen Chakrabarti, Martin van den Berg, and Byron Dom. Focused crawling: a new approach to topic-specific Web resource discovery. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1623–1640, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2178.pdf.

Cohen:1999:LPI

[108] William W. Cohen and Wei Fan. Learning page-independent heuristics for extracting data from Web pages. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1641–1652, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2173.pdf.

Ben-Shaul:1999:ASD

[109] Israel Ben-Shaul, Michael Herscovici, Michal Jacovi, Yoelle S. Maarek, Dan Pelleg, Menachem Shtalhaim, and Vladimir Soroka. Adding support for dynamic and focused search with Fetuccino. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1653–1665, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/

store/comnet/sub/1999/31/11-16/
2171.pdf.

Rossi:1999:IWI

[110] Gustavo Rossi, Daniel Schwabe, and Fernando Lyardet. Improving Web information systems with navigational patterns. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1667–1678, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/11-16/
2141.pdf.

Chakrabarti:1999:SWB

[111] Soumen Chakrabarti, David A. Gibson, and Kevin S. McCurley. Surfing the Web backwards. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1679–1693, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2168.pdf.

Abrams:1999:UAI

[112] Marc Abrams, Constantinos Phanouriou, Alan L. Batongbacal, Stephen M. Williams, and Jonathan E. Shuster. UIML: an appliance-independent XML user interface language. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1695–1708, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1999/31/11-
16/2170.pdf.

Cohen:1999:MTC

[113] Edith Cohen, Haim Kaplan, and Jeffrey Oldham. Managing TCP connections under persistent HTTP. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1709–1723, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/31/11-
                                        16/2144.pdf.

Baker:1999:GNW

[114] Scott Baker and John H. Hartman. The Gecko NFS Web proxy. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1725–1736, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1999/31/11-
                                        16/2165.pdf.

Krishnamurthy:1999:KDB

[115] Balachander Krishnamurthy, Jeffrey C. Mogul, and David M. Kristol. Key differences between HTTP/1.0 and HTTP/1.1. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1737–1751, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/cas/tree/
                                        store/comnet/sub/1999/31/11-16/
                                        2136.pdf.

Cavalli:1999:ATS

[116] Ana Rosa Cavalli. Advanced Topics on SDL and MSC. Computer Networks (Amsterdam, Netherlands: 1999), 31(17):1783–1784, June 7, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/17/2184.pdf.

Mauw:1999:OSM

[117] S. Mauw and M. A. Reniers. Operational Semantics for MSC’96. Computer Networks (Amsterdam, Netherlands: 1999), 31(17):1785–1799, June 7, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/17/2185.pdf.

Mitschele-Thiel:1999:PES

[118] Andreas Mitschele-Thiel and Bruno Müller-Clostermann. Performance engineering of SDL/MSC systems. Computer Networks (Amsterdam, Netherlands: 1999), 31(17):1801–1815, June 7, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/17/2186.
pdf.

Combes:1999:SV

[119] Pierre Combes and Béatrice Renard.

Service validation. Computer Networks (Amsterdam, Netherlands: 1999), 31 (17):1817–1834, June 7, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/17/2187.pdf.

Dssouli:1999:TDC

[120] R. Dssouli, K. Saleh, E. Aboulhamid, A. En-Nouaary, and C. Bourhfir. Test development for communication protocols: towards automation. Computer Networks (Amsterdam, Netherlands: 1999), 31(17):1835–1872, June 7, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/17/2188.
pdf.

Hwang:1999:AGL

[121] Lain-Chyr Hwang and Chung-Ju Chang. Analysis of a general limited scheduling mechanism for a distributed communication system. Computer Networks (Amsterdam, Netherlands: 1999), 31 (18):1879–1889, August 17, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/18/2094.pdf.

Avresky:1999:MDT

[122] D. R. Avresky, V. Shurbanov, R. Wilkinson, R. Horst, W. Watson, and L. Young. Maximum delivery time and hot spots in ServerNet^{T M} topologies. Computer Networks (Amsterdam, Netherlands: 1999), 31(18): 1891–1910, August 17, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/18/2129.pdf.

Leung:1999:PMC

[123] Yiu-Wing Leung. Protocols and minimum capacity for transmission of timecritical message in noisy channel. Computer Networks (Amsterdam, Netherlands: 1999), 31(18):1911–1925, August 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/18/2130.
pdf.

Yaprak:1999:DBA

[124] Ece Yaprak, Anthony Theodore Chronopoulos, Kleanthis Psarris, and Yi Xiao. Dynamic buffer allocation in an ATM switch. Computer Networks (Amsterdam, Netherlands: 1999), 31(18): 1927–1933, August 17, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/18/2132.pdf.

Vandalore:1999:DQC

[125] Bobby Vandalore, Raj Jain, Rohit Goyal, and Sonia Fahmy. Dynamic queue control functions for ATM ABR switch schemes: design and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 31(18): 1935–1949, August 17, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/18/2133.pdf.

Garg:1999:RRR

[126] Rahul Garg and Xiaoqiang Chen. RRR:

recursive round robin scheduler. Computer Networks (Amsterdam, Netherlands: 1999), 31(18):1951–1966, August 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/18/2134.
pdf.

Uyar:1999:TPM

[127] M. Ümit Uyar, Mariusz A. Fecko, Adarshpal S. Sethi, and Paul D. Amer. Testing protocols modeled as FSMs with timing parameters. Computer Networks (Amsterdam, Netherlands: 1999), 31 (18):1967–1988, August 17, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/
comnet/sub/1999/31/18/2135.pdf.

Anonymous:1999:Ea

[128] Anonymous. Editorial(s). Computer Networks (Amsterdam, Netherlands: 1999), 31(19):1995–1997, August 31, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.nl/gej-ng/10/
15/22/32/37/25/abstract.html;
http://www.elsevier.nl/gej-ng/10/
15/22/32/37/25/article.pdf.

Perdikeas:1999:MAS

[129] Menelaos K. Perdikeas, Fotis G. Chatzipapadopoulos, Iakovos S. Venieris, and Gennaro Marino. Mobile agent standards and available platforms. Computer Networks (Amsterdam, Netherlands: 1999), 31(19): 1999–2016, August 31, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/37/26/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/37/26/article.pdf.

Jennings:1999:FCA

[130] Brendan Jennings et al. FIPA-compliant agents for real-time control of Intelligent Network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 31(19):2017–2036, August 31, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/37/27/abstract.html;   http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/37/27/article.pdf.

Breugst:1999:TMS

[131] M. Breugst, L. Faglia, O. Pyrovolakis, I. S. Venieris, and F. Zizza. Towards mobile service agents within an advanced broadband IN environment. Computer Networks (Amsterdam, Netherlands: 1999), 31(19): 2037–2052, August 31, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/37/28/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/37/28/article.pdf.

Lipperts:1999:ABM

[132] Steffen Lipperts and Anthony Sang-Bum Park. An agent-based middleware — a solution for terminal and user mobility. Computer Networks (Amsterdam, Netherlands: 1999), 31(19): 2053–2062, August 31, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/37/29/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/37/29/article.pdf.

Hagen:1999:MAB

[133] L. Hagen, J. Mauersberger, and C. Weckerle. Mobile agent based service subscription and customization using the UMTS virtual home environment. Computer Networks (Amsterdam, Netherlands: 1999), 31(19): 2063–2078, August 31, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/

                                        32/37/30/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/37/30/article.pdf.

Prevedourou:1999:UAT

[134] Didoe Prevedourou et al. Use of agent technology in service and retailer selection in a personal mobility context. Computer Networks (Amsterdam, Netherlands: 1999), 31(19): 2079–2098, August 31, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/37/31/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/37/31/article.pdf.

Mark:1999:FEQ

[135] Jon W. Mark and Jing-Fei Ren. Filtering effects of queueing systems on end-to-end network performance. Computer Networks (Amsterdam, Netherlands: 1999), 31(20): 2103–2113, September 27, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/38/25/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/38/25/article.pdf.

VanMieghem:1999:TIC

[136] Piet Van Mieghem. Topology information condensation in hierarchical networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(20): 2115–2137, September 27, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/38/26/abstract.html;   http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/38/26/article.pdf.

Ma:1999:EMS

[137] Maode Ma, Babak Hamidzadeh, and Mounir Hamdi. An efficient message scheduling algorithm for WDM lightwave networks. Computer Networks (Amsterdam, Netherlands: 1999), 31 (20):2139–2152, September 27, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/38/27/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/38/27/article.pdf.

Bing:1999:MAT

[138] Benny Bing and Regu Subramanian. A multi-access technique for broadband wireless local networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(20):2153–2165, September 27, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/38/28/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/38/28/article.pdf.

Janoska:1999:LAN

[139] Mark W. Janoska and Terence D.

Todd. Local area networks using a spatial wavelength cross connect. Computer Networks (Amsterdam, Netherlands: 1999), 31(20):2167– 2185, September 27, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/38/29/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/38/29/article.pdf.

Anonymous:1999:CP

[140] Anonymous. Conference paper(s). Computer Networks (Amsterdam, Netherlands: 1999), 31(20):2189–2191, September 27, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/15/22/32/38/30/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/15/22/32/38/30/article.pdf.

Sidorovs:1999:SMI

[141] Andris Sidorovs, Guntis Barzdins, Janis Lacis, and Karlis Ogsts. SmartARP: merging IP and MAC addressing for low-cost gigabit Ethernet networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(21): 2193–2204, November 10, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/26/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/26/article.pdf.

Tully:1999:LFW

[142] John Tully and Arnis Riekstins. Licencefree wireless internet access technologies. Computer Networks (Amsterdam, Netherlands: 1999), 31(21): 2205–2213, November 10, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/27/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/27/article.pdf.

Olbrich:1999:VRM

[143] S. Olbrich and H. Pralle. Virtual reality movies-real-time streaming of 3D objects. Computer Networks (Amsterdam, Netherlands: 1999), 31(21): 2215–2225, November 10, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/40/28/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/40/28/article.pdf.

Jarkowski:1999:NAM

[144] Maciej Jarkowski, Artur Binczewski, and Maciej Stroinski. New approach for management services with a Web browser. Computer Networks (Amsterdam, Netherlands: 1999), 31(21): 2227–2236, November 10, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/40/29/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/40/29/article.pdf.

Nentwig:1999:LVE

[145] Lutz Nentwig. LeMO: A virtual exhibition of 20th century German history. Computer Networks (Amsterdam, Netherlands: 1999), 31(21): 2237–2244, November 10, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/40/30/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/40/30/article.pdf.

Berg:1999:CEM

[146] Alexander von Berg and Helmut Pralle.

A concept for an electronic magazine. Computer Networks (Amsterdam, Netherlands: 1999), 31(21): 2245–2251, November 10, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/31/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/31/article.pdf.

Sabatino:1999:IPE

[147] Roberto Sabatino and Jose Manuel de Arce. Implementation of the pan-European academic research network: TEN-155. Computer Networks (Amsterdam, Netherlands: 1999), 31(21): 2253–2261, November 10, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/32/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/32/article.pdf.

Overmeer:1999:SIM

[148] Mark A. C. J. Overmeer. A search interface for my questions. Computer Networks (Amsterdam, Netherlands: 1999), 31(21):2263–2270, November 10, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/33/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/33/article.pdf.

Overmeer:1999:MPS

[149] Mark A. C. J. Overmeer. My personal search engine. Computer Networks (Amsterdam, Netherlands: 1999), 31(21): 2271–2279, November 10, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/34/abstract.html;   http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/40/34/article.pdf.

Palme:1999:TBW

[150] Jacob Palme. Talking back to the WWW. Computer Networks (Amsterdam, Netherlands: 1999), 31(21): 2281–2286, November 10, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/40/35/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/40/35/article.pdf.

Lopez:1999:PSM

[151] Diego R. López and Marcelo Reina.

Providing secure mobile access to information servers with temporary certificates. Computer Networks (Amsterdam, Netherlands: 1999), 31(21): 2287–2292, November 10, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/40/36/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/40/36/article.pdf.

Lizcano:1999:MSA

[152] Pedro J. Lizcano, Arturo Azcorra, Josep Solé-Pareta, Jordi Domingo-Pascual, and Manuel Alvarez-Campana. MEHARI: a system for analysing the use of the internet services. Computer Networks (Amsterdam, Netherlands: 1999), 31(21):2293–2307, November 10, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/40/37/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/40/37/article.pdf.

Niederberger:1999:HSS

[153] Ralph Niederberger, Helmut Grund, Eva Pless, and Ferdinand Hommes. High-speed supercomputer communications in broadband networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(21):2309–2318, November 10, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.nl/gej-ng/10/
                                        15/22/32/40/38/abstract.html;
                                        http://www.elsevier.nl/gej-ng/10/
                                        15/22/32/40/38/article.pdf.

Anonymous:1999:Eb

[154] Anonymous. Editorial(s). Computer Networks (Amsterdam, Netherlands: 1999), 31(21):iii, November 10, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/25/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/40/25/article.pdf.

DeMartino:1999:II

[155] Kevin DeMartino. ISDN and the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 31(22): 2325–2339, November 29, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/41/25/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/41/25/article.pdf.

Cobb:1999:UTS

[156] Jorge A. Cobb. Universal Timestamp-Scheduling for real-time networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(22):2341–2360, November 29, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/15/22/32/41/26/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/15/22/32/41/26/article.pdf.

Hassan:1999:HRR

[157] Mahbub Hassan, James Breen, and Mohammed Atiquzzaman. HG-RCP: A rate-based flow control mechanism for intranets interconnected by ATM networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(22): 2361–2379, November 29, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/41/27/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/41/27/article.pdf.

Naldi:1999:MBM

[158] M. Naldi. Measurement-based modelling of Internet dial-up access connections. Computer Networks (Amsterdam, Netherlands: 1999), 31(22): 2381–2390, November 29, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/41/28/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/41/28/article.pdf.

Kamal:1999:DPW

[159] A. E. Kamal and G. K. Janssens. Design, performance and wavelength assignment of a wavelength division multiaccess protocol for optical fibre ring networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(22): 2391–2410, November 29, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/41/29/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/41/29/article.pdf.

Yau:1999:OWN

[160] Victor Yau. Optical WDMA network design and function placement. Computer Networks (Amsterdam, Netherlands: 1999), 31(22): 2411–2427, November 29, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/41/30/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/41/30/article.pdf.

Dacier:1999:ID

[161] Marc Dacier and Kathleen Jackson. Intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 31 (23–24):2433–2434, December 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/42/25/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/42/25/article.pdf.

Paxson:1999:BSD

[162] Vern Paxson. Bro: a system for detecting network intruders in realtime. Computer Networks (Amsterdam, Netherlands: 1999), 31(23–24): 2435–2463, December 14, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/42/26/abstract.html;   http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/42/26/article.pdf.

Huang:1999:LSD

[163] Ming-Yuh Huang, Robert J. Jasper, and Thomas M. Wicks. A large scale distributed intrusion detection framework based on attack strategy analysis. Computer Networks (Amsterdam, Netherlands: 1999), 31(23–24): 2465–2475, December 14, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/42/27/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/42/27/article.pdf.

Sommer:1999:IDS

[164] Peter Sommer. Intrusion detection systems as evidence. Computer Networks (Amsterdam, Netherlands: 1999), 31 (23–24):2477–2487, December 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/42/28/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/42/28/article.pdf.

Sharon:1999:SMB

[165] Oran Sharon and Adrian Segall. Session management in broadcast networks with dynamic addresses. Computer Networks (Amsterdam, Netherlands: 1999), 31 (23–24):2489–2503, December 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/42/29/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/42/29/article.pdf.

Cankaya:1999:ISA

[166] Hakki C. Cankaya and V. S. S. Nair. Improved survivability analysis for SONET SHRs. Computer Networks (Amsterdam, Netherlands: 1999), 31(23–24): 2505–2528, December 14, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/42/30/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/42/30/article.pdf.

Carter:1999:NID

[167] Robert L. Carter and Mark E. Crovella. On the network impact of dynamic server selection. Computer Networks (Amsterdam, Netherlands: 1999), 31 (23–24):2529–2558, December 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/42/31/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/15/22/
                                        32/42/31/article.pdf.

Peha:1999:SAC

[168] Jon M. Peha. Scheduling and admission control for integrated-services networks: the Priority Token Bank. Computer Networks (Amsterdam, Netherlands: 1999), 31(23–24):2559–2576, December 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.nl/gej-ng/10/
                                        15/22/32/42/32/abstract.html;
                                        http://www.elsevier.nl/gej-ng/10/
                                        15/22/32/42/32/article.pdf.

Kolberg:1999:EMU

[169] Mario Kolberg, Richard O. Sinnott, and Evan H. Magill. Experiences modelling and using formal objectoriented telecommunication service frameworks. Computer Networks (Amsterdam, Netherlands: 1999), 31(23–24): 2577–2592, December 14, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/42/33/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/42/33/article.pdf.

Anonymous:1999:Ia

[170] Anonymous. Index. Computer Networks (Amsterdam, Netherlands: 1999), 31 (23–24):2597–2605, December 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/42/34/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/42/34/article.pdf.

Anonymous:1999:Ib

[171] Anonymous. Index. Computer Networks (Amsterdam, Netherlands: 1999), 31 (23–24):2607–2612, December 14, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/15/22/
32/42/35/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/15/22/
32/42/35/article.pdf.

Albrecht:1999:ISB

[172] M. Albrecht, M. Frank, P. Martini, and M. Schetelig. IP services over Bluetooth: Leading the way to a new mobility. In IEEE [207], pages 2–13. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

VanHoudt:1999:PLP

[173] B. Van Houdt, C. Blondia, O. Casals, and J. Garcia. Packet level performance characteristics of a MAC protocol for wireless ATM LANs. In IEEE [207], pages 14–23. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Coutras:1999:EPH

[174] C. Coutras and P.-J. Wan. Evaluating performance of the HIPERLAN CAC layer protocol for asynchronous traffic. In IEEE [207], pages 24–33. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Bing:1999:MPI

[175] B. Bing. Measured performance of the IEEE 802.11 wireless LAN. In IEEE [207], pages 34–42. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Georgantas:1999:MBM

[176] N. Georgantas, N. Papadoukakis, A. Liveris, and E. Sykas. Modelling of a bandwidth management scheme for DECT wireless access to a local area network. In IEEE [207], pages 43–48. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Hope:1999:DPR

[177] M. Hope and N. Linge. Determining the propagation range of IEEE 802.11 radio LANs for outdoor applications. In IEEE [207], pages 49–51. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Matijasevic:1999:IMN

[178] M. Matijasevic, D. Gracanin, K. Valavanis, and I. Lovrek. Interconnection model for networked virtual reality applications. In IEEE [207], pages 52–61. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Takahara:1999:VBE

[179] A. Takahara, S.-I. Tani, S. Ishihara, and T. Miyazaki. Virtual BUS: An easyto-use environment for distributed resources. In IEEE [207], pages 62–70. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Morlang:1999:PNS

[180] A. Morlang and I. Schieferdecker. On programming of network services. In IEEE [207], pages 71–72. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Roy:1999:DGA

[181] R. Roy. Distributed gatekeeper architecture for H.323-based multimedia telephony. In IEEE [207], pages 73–76. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Jha:1999:SSQ

[182] S. Jha and A. Seneviratne. Synchronization skew: A QoS measurement study. In IEEE [207], pages 77–79. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Salah:1999:PRO

[183] K. Salah, E. Drakopoulos, and T. Elrad. Periodic route optimization for handed-off connections in wireless ATM networks. In IEEE [207], pages 80–87. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Vandalore:1999:AAQ

[184] B. Vandalore, R. Jain, S. Fahmy, and S. Dixit. AQuaFWiN: Adaptive QoS framework for multimedia in wireless networks and its comparison with other QoS frameworks. In IEEE [207], pages 88–97. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Lazar:1999:AND

[185] S. Lazar, D. Sidbu, S. Kodeswaran, and R. Varadharaj. ATM network discovery using mobile agents. In IEEE [207], pages 98–105. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Shankaran:1999:SME

[186] R. Shankaran, V. Varadharajan, and M. Hitchens. Secure multicast extensions for mobile networks. In IEEE [207], pages 106–116. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Atwood:1999:UFS

[187] J. Atwood, M. Ghodrat, and D. Tasak.

Using formal specification and observers to specify and validate the ATM signaling protocols. In IEEE [207], pages 117– 121. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Guenter:1999:AMQ

[188] M. Guenter, T. Braun, and I. Khalil. An architecture for managing QoS-enabled VPNs over the Internet. In IEEE [207], pages 122–131. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Varada:1999:DFB

[189] S. Varada, V. Oduol, and A. Shelat. Data flow and buffer management in multi-channel data link controller. In IEEE [207], pages 132–141. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Wan:1999:ORB

[190] P.-J. Wan, L. Liu, and Y. Yang. Optimal routing based on super topology in hypercube WDM networks. In IEEE [207], pages 142–149. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Aboelela:1999:FTR

[191] E. Aboelela and C. Douligeris. Fuzzy temporal reasoning model for event correlation in network management. In IEEE [207], pages 150–159. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Feuser:1999:EIF

[192] O. Feuser and A. Wenzel. On the effects of the IEEE 802.3x flow control in full-duplex Ethernet LANs. In IEEE [207], pages 160–163. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

You:1999:TSM

[193] C. You and K. Chandra. Time series models for Internet data traffic. In IEEE [207], pages 164–171. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Elteto:1999:DRT

[194] T. Elteto and S. Molnar. On the distribution of round-trip delays in TCP/IP networks. In IEEE [207], pages 172–181. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Wang:1999:EAE

[195] J. Wang and S. Keshav. Efficient and accurate Ethernet simulation. In IEEE [207], pages 182–191. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Zhang:1999:WCS

[196] S. Zhang and E. Lee. The worst-case scenario for transmission of synchronous traffic in an FDDI network. In IEEE [207], pages 192–203. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Fujinoki:1999:NSB

[197] H. Fujinoki and K. Christensen. The new shortest best path tree (SBPT) algorithm for dynamic multicast trees. In IEEE [207], pages 204–211. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Baatz:1999:FAS

[198] S. Baatz, M. Frank, P. Martini, and A. Wenzel. Flow aggregation and scalability with ROBIN: End-to-end regulation of bandwidth in intranetworks. In IEEE [207], pages 212–223. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Pinto:1999:APP

[199] J. Pinto and K. Christensen. An algorithm for playout of packet voice based on adaptive adjustment of talkspurt silence periods. In IEEE [207], pages 224– 231. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Ishibashi:1999:MSS

[200] Y. Ishibashi, S. Tasaka, and Y. Tachibana.

A media synchronization scheme with causality control in network environments. In IEEE [207], pages 232–241. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Amanton:1999:CPO

[201] L. Amanton and M. Naimi. A causal-phase ordering protocol with multi-initiators for overlapped broadcasts. In IEEE [207], pages 242–253. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Gilfix:1999:ISI

[202] M. Gilfix. An integrated software immune system: A framework for automated network management, system health, and security. In IEEE [207], pages 254–255. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Choi:1999:DBD

[203] W.-C. Choi. Decoupling bandwidth, delay and burstiness of PFQ algorithm for QoS guarantee of a leaky bucket constraint traffic. In IEEE [207], pages 256– 257. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Chang:1999:NPS

[204] W.-C. Chang, C.-H. Ko, Y.-H. Lee, and S.-T. Sheu. A novel prediction system for wireless LAN based on the genetic algorithm and neural network. In IEEE [207], pages 258–259. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Callahan:1999:IHP

[205] D. Callahan and G. Grimes. An intelligent hub protocol for local area lightwave networks. In IEEE [207], pages 260–261. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.

Anonymous:1999:PEI

[206] Anonymous, editor. Proceedings of the Eighth International World Wide Web Conference: Toronto, Canada, 11–14 May 1999, volume 31(11–16) of Computer Networks and ISDN Systems. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1999. ISBN 0-444-50264-5. LCCN TK5105.875.I57 P76 1999.

IEEE:1999:LPC

[207] IEEE, editor. LCN’99: proceedings:

24th Conference on Local Computer Networks: October 18–20, 1999, Lowell, Massachusetts, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7695-0309-8, 0-7695-0311-X (microfiche). LCCN TK5105.5 .C66 1999. IEEE Computer Society Order Number PR00309.